OpenStudy (anonymous):

elimination 3x-2y=11 3x-y=7

OpenStudy (texaschic101):

3x - 2y = 11 3x - y = 7 ---->(-1) ---------------- 3x - 2y = 11 -3x + y = -7 (result of multiplying by -1) ----------------add -y = 4 -- multiply by -1 to make y positive y = -4 now sub -4 in for y in either equation to find x 3x - 2y = 11 3x - 2(-4) = 11 3x + 8 = 11 3x = 11 - 8 3x = 3 x = 1 check... 3x - y = 7 3(1) - (-4) = 7 3 + 4 = 7 7 = 7 (correct) x = 1 and y = -4 any questions ?